OpenStudy (anonymous):

rubidium changes from a soild to a liquid at 102.74f. what is this temperature in celsisus? use the formula c=5^9(f-32) to find the celsius temperature.

OpenStudy (whpalmer4):

The formula is \[C = \frac{5}{9}(F-32)\] \[C = \frac{5}{9}(102.74-32) = \frac{5}{9}(70.74) = \]